# What does really matter in image goal navigation?

### v1

@@ 5 Discussion and conclusion
In this paper we have studied the connections between image goal navigation and relative pose estimation. We have shown that the success of recent late fusion architectures based on channel stacking can be linked to a simulator setting allowing the agent to slide along walls. While agents trained with sliding enabled are known to transfer badly to real environments, where this is not possible, our analysis has shown that capabilities learned with this sliding setting can actually be partially transferred to more realistic environment settings. While sliding is a property of agent dynamics and not inherently linked to perception, we have shown that it significantly impacts training of the visual encoders of end-to-end trained agents: transferring capabilities to the realistic setting only seems to be successful if it also involves transferring the visual encoder weights. We conjecture that training with the realistic setting (sliding disabled) leads to undertraining both the perception module, as the goal is seen more rarely, as well as the rest of the agent, which has a difficulty to cover later portions of the episodes.
We also compared and analyzed different widely-used architectures: Late Fusion, ChannelCat, SpaceToDepth projections and Cross-attention. We conjecture that successful architectures require support for early fusion of representations, allowing to compute correspondences between local parts of images, and we found evidence for this in our experiments: Late Fusion generally underperforms. As expected, our experiments also found a correlation between navigation performance and (emerging) relative pose estimation performance, which we probed with explicit heads. Finally, we argue that, up to now and up to our knowledge, there does not seem to exist a simple solution to learn image goal navigation successfully end-to-end with simple relatively low-capacity architectures and without pre-training, which we judge to remain essential for goal oriented navigation.
Appendix A Implementation details
ResNet backbone – The ResNet implementation is borrowed from [56, 1, 67] and the code is publicly available  [^7] https://github.com/XinyuSun/FGPrompt. The default network is implemented as a custom ResNet9 with SpaceToDepth blocks [56] that rearrange the input of each 2D convolution layer such that each of its spatial dimensions is divided by a factor 4, while the channel dimension is 16 times larger. We consider a ResNet9 because of the performance reported in [67], in line with those of a ResNet50, but with a fraction of the parameters. In the ResNet version without SpaceToDepth, these reshaping blocks are simply skipped. As customary, the 2D CNN output is flatten and passed to a linear layer to generate a 128D feature vector.
@@ Appendix A Implementation details
ViT backbone – We use the small version of a popular ViT implementation [^8] https://github.com/huggingface/pytorch-image-models, augmented with the convolutional compression layer proposed in OVRL-v2 [75] (code available in the paper) to compress the ViT output to a 2058D feature. As discussed in [75], without pre-training (as is the case here), ViT-Small performs better than Base despite having a fraction of the parameters, we thus use ViT-Small for all our experiments. We adopt this layer since Yadav et al. [75] report significant improvements in navigation performance using it, over more common compression options like average pooling or [CLS] token.
@@
DEBiT backbone – In the paper we use the DEBiT-b version of the architecture, the variant of DEBiT with the best performance-size trade-off and publicly available with its pre-trained binocular encoder  [^9] https://github.com/naver/debit. The binocular encoder is kept frozen and is based on the Geometric Foundation Model CroCo [71], finetuned to estimate relative pose and visibility as described in [8] and the main text. The network is implemented with two Siamese encoders, applied to observation and goal images, and a decoder which combines the output of these two encoders. Both encoders and decoders are implemented with a ViT architecture with self-attention layers, with the decoder also featuring cross-attention. The output of the decoder is further compressed by a fully connected layer (also kept frozen, from the DEBiT model) that projects the flattened output of the decoder into a 3136D feature vector.
@@
Table 6: Comparisons of DEBiT [8] in two settings: our standard 2-layer GRU test architecture used in the paper (Tables 1,2 and 3), and the architecture originally reported in [8] (1-layer GRU).
Appendix B Additional experiments
Impact of GRU structure – For fairness of comparison, we used the same GRU architecture in all our experiments, with 2 layers and hidden state of size 128. In Table 6 we show the results for training DEBiT-b with the same GRU architecture which reported in [8], with one layer and hidden dimension 512. The experiment excludes any dependency on this choice: the results are very similar in spite of the different number of layers and the change in sizes of the recurrent memory $|\mathbf{h}_{t}|$.

### v3

@@ 5 Discussion and conclusion
We have studied image goal navigation and raised questions on agent architecture and whether components dedicated to vision or decision taking should be trained jointly with RL or separately. We have shown that the success of recent architectures based on channel stacking can be linked to a simulator setting allowing sliding. While agents trained with sliding enabled are known to transfer badly to real environments, where this is not possible, our analysis has shown that capabilities learned with this sliding setting can actually be partially transferred to more realistic environment settings. While sliding is a property of agent dynamics and not inherently linked to perception, we have shown that it significantly impacts training of the visual encoders of end-to-end trained agents: transferring capabilities to the realistic setting only seems to be successful if it also involves transferring the visual encoder weights. We conjecture that training with the realistic setting leads to undertraining both the perception module, as the goal is seen more rarely, and the rest of the agent, which has a difficulty to cover later portions of the episodes.
We also compared and analyzed different widely-used architectures: Late Fusion, ChannelCat, SpaceToDepth projections and Cross-attention. Results suggest that successful architectures require support for early fusion of representations, allowing to compute correspondences between local parts of images: Late Fusion generally underperforms. As expected, we also found a correlation between navigation performance and (emerging) relative pose estimation performance, which we probed with explicit heads. Finally, we argue that, up to our knowledge, there does not seem to exist a simple solution to learn with RL alone image goal navigation end-to-end with simple low-capacity architectures and without pre-training, which we judge to remain essential for goal-oriented navigation.
Appendix A Agents implementation details
ResNet backbone – The ResNet implementation is borrowed from [60, 1, 70] and the code is publicly available. [^6] https://github.com/XinyuSun/FGPrompt The default network is implemented as a custom ResNet9 with SpaceToDepth blocks [60] that rearrange the input of each 2D convolution layer such that each of its spatial dimensions is divided by a factor 4, while the channel dimension is 16 times larger. In the paper we use the ResNet9 version because of the performance reported in [70], in line with those of a ResNet50, but with a fraction of the parameters. In the ResNet version without SpaceToDepth, these reshaping blocks are simply skipped. As customary, the 2D CNN output is flatten and passed to a linear layer to generate a 128D feature vector.
@@ Appendix A Agents implementation details
ViT backbone – We use the small version of a popular ViT implementation, [^7] https://github.com/huggingface/pytorch-image-models augmented with the convolutional compression layer proposed in OVRL-v2 [78] (code available in the Appendix of the OVRL-v2 paper) to compress the ViT output to a 2058D feature. We adopt this layer since [78] reports significant improvements in navigation performance using it, over more common compression options like average pooling or [CLS] token. As discussed in [78], without pre-training (as is the case here), ViT-Small performs better than ViT-Base despite having a fraction of the parameters, and therefore we use ViT-Small for all our experiments in the paper.
@@
DEBiT backbone – In the paper we use the DEBiT-B version of the architecture, the variant of DEBiT with the best performance-size trade-off and publicly available with its pre-trained binocular encoder. [^8] https://github.com/naver/debit The binocular encoder is kept frozen and is based on the Geometric Foundation Model CroCo [74], finetuned to estimate relative pose and visibility as described in [8] and the main text. The network is implemented with a Siamese encoder, applied to observation and goal images, and a decoder which combines the output of these two encoders. Both encoders and decoders are implemented with a ViT architecture with self-attention layers, with the decoder also featuring cross-attention. The output of the decoder is further compressed by a fully connected layer (also kept frozen, from the DEBiT model) that projects the flattened output of the decoder into a 3136D feature vector.
@@
We use the same DEBiT architecture for the variant without pre-trained binocular encoder, the only difference being that the binocular encoder is trained from scratch with the other components of the agent using only RL. Similarly to what is reported in [8], we were not able to train such large architecture to a usable state using RL alone.
Model ZSEL [1]
Repository https://github.com/ziadalh/zero_experience_required
Configs config/imagenav/eval_ppo_imagenav_rgb.yaml
$\hookrightarrow$ config/imagenav/gibson/imagenav_rgb.yaml
Link github.com/ziadalh/zero_experience_required/blob/main/config/imagenav/gibson/imagenav_rgb.yaml#L15
Model ZSON [46]
Repository https://github.com/gunagg/zson
Configs configs/tasks/imagenav_hm3d.yaml
Link github.com/gunagg/zson/blob/main/configs/tasks/imagenav_hm3d.yaml#L13
Model PIRLNav [56]
Repository https://github.com/Ram81/pirlnav
Configs configs/experiments/il_objectnav.yaml
$\hookrightarrow$ configs/tasks/objectnav_hm3d.yaml
Link github.com/Ram81/pirlnav/blob/main/configs/tasks/objectnav_hm3d.yaml#L15
Model GOAT-Bench [35]
Repository https://github.com/Ram81/goat-bench
Configs config/experiments/ver_goat_skill_chain.yaml
$\hookrightarrow$ config/tasks/goat_stretch_hm3d.yaml
Link github.com/Ram81/goat-bench/blob/main/config/tasks/goat_stretch_hm3d.yaml#L78
Model DEBiT [8]
Repository https://github.com/naver/debit
Configs configs/imgnav-gibson-debit.yaml
Link github.com/naver/debit/blob/main/configs/imgnav-gibson-debit.yaml#L15
Model FGPrompt [70]
Repository https://github.com/XinyuSun/FGPrompt
Configs exp_config/ddppo_imagenav_gibson.yaml
$\hookrightarrow$ exp_config/imagenav_gibson.yaml
Link N/A
Table 5: How the allow_sliding flag is set in other codebases. Here we list, for each method, the repository address, the main configuration file used for the main experiment (with $\hookrightarrow$ we indicate the eventual config file composed with the main one where the Sliding flag is set), and the link to the line where the parameter is set. FGPrompt does not appear to set Sliding to False, thus it uses the default value of True.
Appendix B Sliding setting in Habitat
As discussed in the main paper, the allow_sliding parameter, or Sliding in short, in the Habitat simulator impacts the way an agent behaves when colliding with obstacles: if set to True the agent slides along obstacles, while when False the agent stops. By default the flag is switched to True, but Kadian et al. [33] showed that setting Sliding=True has a big detrimental impact in sim2real transfer. Therefore, it is common practice in the field to set Sliding=False to train and test agents in conditions that are as realistic as possible.
We analyzed several publicly-available code repositories to assess how these set the Sliding flag. As already mentioned, by default it is True, and it is usually modified in a yaml configuration file. Habitat uses Hydra [^9] https://hydra.cc/ to manage the configuration, which allows to compose multiple configuration files, making it cumbersome sometimes to track the exact final configuration used for each experiment. In Tab. 5 we list, for several relevant robot navigation methods using Habitat, the link to the repository, the configuration file used for the main experiment (and other config files composed with the main one indicated by $\hookrightarrow$), and a link to the line in the code where allow_sliding is set to False. Furthermore, we have searched the repositories for occurrences of the string “allow_sliding” to verify that it is not modified directly in the code. All methods but FGPrompt set Sliding to False, hence this codebase uses the default setting of allow_sliding=True. For completeness, we note that we could find an occurrence of allow_sliding=False in a config file in the FGPrompt repository, [^10] https://github.com/XinyuSun/FGPrompt/blob/main/src/config/v2.yaml#L11 but this file does not appear to be used anywhere in the code. As a final check, we installed and ran the code of FGPrompt with the recommended settings and commands, and we could indeed verify that allow_sliding=True in the Habitat environments used for training and testing.
Appendix C Additional experiments
Impact of GRU structure – For fairness of comparison, we used the same GRU architecture in all our experiments, with 2 layers and hidden state of size 128. In Table 6 we show the results for training DEBiT-B with the same GRU architecture reported in [8], with one layer and hidden dimension 512. The experiment excludes any dependency on this choice: the results are very similar in spite of the different number of layers and the change in sizes of the recurrent memory $|\mathbf{h}_{t}|$.
@@ Appendix C Additional experiments
Table 6: Comparisons of DEBiT [8] in two settings: our standard 2-layer GRU test architecture used in the paper (Table 1), and the architecture originally reported in [8] (1-layer GRU).